OpenStudy (anonymous):

WILL GIVE MEDALS!!! Is x - 8 a factor of the function f(x) = -2x³ + 17x² - 64? Explain

OpenStudy (anonymous):

find f(8) if it is 0 then x-8 is a factor.

OpenStudy (anonymous):

oh okay!

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

@INeedSomeHelp!

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

you need to do long division

OpenStudy (anonymous):

no it is not because there is a remainder of 8

OpenStudy (anonymous):

\[f(8)=-2(8)^3+17(8)^2-64=8^2\left( -2*8+17-1 \right)=64*0=0\] hence x-8 is a factor of f(x)

OpenStudy (anonymous):

hmm okay! thanks everyone!!!